OpenStudy (anonymous):

I understand that this may be a pretty lame question, but it's been a big pain to me all day. Does anyone know how I can get the least common denominator to add these two fractions? (1/5.1) + (1/12) thank you in advanced!

OpenStudy (anonymous):

1/5.1 can be reduced to 10/51

OpenStudy (anonymous):

now you can find the LCM of 51 and 12 51=3*17 12=2*2*3 Lcm of 51 and 12 is given by 2*2*3*17= 204

OpenStudy (anonymous):

Ohmg, thank you so much for your help. I really appreciate it!!

OpenStudy (anonymous):

yw